If c∈Vn, show that the function f given by f(x)=c.x is continuous on R^n.
1
Expert's answer
2020-11-27T02:57:00-0500
"Solution"

The function "f: \rm I\!R^2 \to \rm I\!R" defined by "f(x)=c \cdot x" for a fixed "c".


I'll show that "f" is continuous at a point "x_0 \in \rm I\!R^n" . Given any "e >0", we need to show that there exists "\delta >0" such that "\| x-x_0\|<\delta \implies |f(x)-f(x_0)|<e"


Now consider "|f(x)-f(x_0)|=|c \cdot x-c \cdot x_0|=|c \cdot (x-x_0)| \leq \|c\|\ \|x-x_0\|"


Choose "\delta= \frac{\|c\|}{e} \implies |f(x)-f(x_0)|\leq \|c\| \|x-x_0\|<\|c\| \delta=e"